OpenStudy (princeevee):

simplify -9+(-5)x+3+(-2)x -12-7x -6+3x -6-7x 3-16x

Vocaloid (vocaloid):

combine like terms

Vocaloid (vocaloid):

-5x - 2x = ?

OpenStudy (princeevee):

so its c?

OpenStudy (princeevee):

thanks one more?

OpenStudy (princeevee):

1 -(a+b)+a=-a+(-b)+A statement: distributive property 2 =a+a+(-b) statement communitive property of addition 3 =0+(-B) statement: additive inverse property = - b statement: ? what is the statement for 4?

OpenStudy (princeevee):

@tracyplays

OpenStudy (princeevee):

?

OpenStudy (princeevee):

no thats it

OpenStudy (princeevee):

what is the reason for it

OpenStudy (princeevee):

is it additive inverse definition of subtraction commutative property of addition or identity property of addition?